A snail travelling 1 m/hr and speeds up to 7m/hr in 2hrs. What is the acceleration?


1
Expert's answer
2019-09-19T09:48:19-0400

Let's start from the initial acceleration formula:

"a = (V_2 - V_1)/\Delta t"

Find the acceleration of the snail:

"a = (7 - 1)\/2 = 3 (m\/hr^2)"

Answer: 3 m/hr2 -- the acceleratin of the snail.